OpenStudy (anonymous):

Janet told Sam that she has read 10 books this year and reads 2 books each month. Sam wants to catch up to Janet. He tracks his book reading with a table on his door. Using his table below, what month will Sam have read the same amount of books as Janet?

OpenStudy (anonymous):

MonthBooks June 3 July 6 August 9 @phi

OpenStudy (phi):

they want you to make up the equation for two different lines slope is "books read per month" x is month if we assign numbers to the months June=6, July=7, then you can change the table to 6,3 7,6 8,9 you have 3 points. You can use any two to find the equation of the line y is the number of books read find the equation, then set y =10 and solve for x (the month)

OpenStudy (anonymous):

im confused @phi

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@thomaster

thomaster (thomaster):

June 3 July 6 August 9 How many books does sam read each month?

OpenStudy (anonymous):

3

thomaster (thomaster):

Janet reads 2 books each month So in June, she had read 12 books Sam has read 3 books in June For July, it's 14 for janet and 6 for Sam August 16 and 9 September 18 and 12 November 20 and 15 Now continue until they have both read the same amount of books.

OpenStudy (anonymous):

ok 22 18 24 21 26 24 28 27 30 30

OpenStudy (anonymous):

November December January February

thomaster (thomaster):

So in which month did they both reached the total of 30 books?

OpenStudy (anonymous):

november

OpenStudy (anonymous):

What equation is solved by the graphed systems of equations?

OpenStudy (anonymous):

3x + 4 = 5x - 10 3x - 4 = 5x - 10 3x + 4 = 5x + 10 3x - 4 = 5x + 10

thomaster (thomaster):

november is wrong

thomaster (thomaster):

20 15 = november 22 18 24 21 26 24 28 27 30 30

OpenStudy (anonymous):

then how do i find it

thomaster (thomaster):

20 15 = november 22 18 = december 24 21 = january 26 24 28 27 30 30

OpenStudy (anonymous):

Janet told Sam that she has read 10 books this year and reads 2 books each month. Sam wants to catch up to Janet. He tracks his book reading with a table on his door. Using his table below, what month will Sam have read the same amount of books as Janet? --- 10 books this year 2 book each month = 24 per year express it like this 2x+10 = total books read Janet where x is month now for sam ax+b=sam total books. Where x is month a is howmany per month and b is how many books has he read this year so far.

OpenStudy (anonymous):

you can make june =1 for x and july 2 and so on and so on.

OpenStudy (anonymous):

i wanna say january